An object initially traveling at 30. m/s south accelerates uniformly at 5.0 m/s/s north and is displaced 10 meters. The final velocity of the object is....


1
Expert's answer
2021-01-14T12:51:54-0500

The velocity of an object under a constant acceleration is given as follows:


"v = v_0 - at"

where "v_0 = 30m\/s" is the initial speed, "a= 5m\/s^2" is the acceleration, "t" is the time of motion. Sign minus here is due to the opposite directions of the initial velocity and acceleration (south and north respectively).

The distance "d = 10m" travelled by the object is given as follows:


"d = v_0t - \dfrac{at^2}{2}"

Solving the quadratic equation for "t", obtain:


"t = \dfrac{v_0 + \sqrt{v_0^2 - 2ad}}{a}"

Substituting into the first equation, obtain:


"v = v_0 - v_0-\sqrt{v_0^2 - 2ad} = -\sqrt{v_0^2 - 2ad}\\
v = -\sqrt{30^2 - 2\cdot 5\cdot 10} \approx -28.3m/s"

Answer. -28.3 m/s.